OpenStudy (anonymous):

What is the solution to 2x2 + x + 2 = 0?

OpenStudy (anonymous):

@agent0smith

OpenStudy (anonymous):

Is that 2 times 2?

OpenStudy (anonymous):

its supposed to be 2x squared

OpenStudy (anonymous):



OpenStudy (anonymous):

\[\large 2x^2+x+2\]

OpenStudy (anonymous):

^ like that lol

OpenStudy (anonymous):

\[\large \frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\]

OpenStudy (anonymous):

recognise that formula?

OpenStudy (anonymous):

or did they not teach you that yet?

OpenStudy (anonymous):

i recognize it i just am not real sure how to complete it once i plug it in

OpenStudy (anonymous):

alright, which part of it are you unsure about than?

OpenStudy (anonymous):

all of it...im so behind on my work so nothing is making sense because im so overwhelmed.

OpenStudy (anonymous):

i know the feeling.

OpenStudy (anonymous):

anyway. the formula above holds try for all equations in this format: \[ax^2+bx+c=0\]

OpenStudy (anonymous):

your equation is exactly like that. so it's pretty much a 1-on-1 copy.

OpenStudy (anonymous):

the number in front of the x² is your a, the number in front of your x is your b (if there is no number in front of x you must use 1 instead, if there is no x at all then use 0), the number without an x is your c.

OpenStudy (anonymous):

does that make sense?

OpenStudy (anonymous):

if that makes sense to you, please tell me which numbers you would pick for a, b and c.

OpenStudy (anonymous):

2,1,2?

OpenStudy (anonymous):

indeed.

OpenStudy (anonymous):

i'll fill them in for you

OpenStudy (anonymous):

\[\frac{ -1\pm \sqrt{1^2-4*2*2}}{ 2*2 }\]

OpenStudy (anonymous):

solving that will give you the values of x. In this case you can't solve it though, the number under the root is negative.

OpenStudy (anonymous):

\[\frac{ -1\sqrt{1^{2}}-16 }{ 4 }\]

OpenStudy (anonymous):

thats what i have got so far idk how to do it from here though

OpenStudy (anonymous):

did you accidentally miss the 16 under the root or is it supposed to be like that?

OpenStudy (anonymous):

if you want to have multiple numbers under a root you can write {} after the root command anything between those two will be under the root.

OpenStudy (anonymous):

oh ok, well yes it was supposed to go under it

OpenStudy (anonymous):

ye, you can't solve that, means x has no solution. (if you draw a graph you can prove it because x does not actually come close to 0, it doesn't go below 2.

OpenStudy (anonymous):